# Neutral-current background induced by atmospheric neutrinos at large liquid-scintillator detectors: III. Quantitative calculations for reactor neutrinos

### v1

@@ 2.3 GEANT4-based Detector Simulation
GEANT4 (version 4.10.p02) is utilized to simulate the propagation of final-state particles resulting from -C NC interactions within the LS detector. The simulation employs the QGSPBERT physics list, with the Quark Gluon String (QGS) model applied for hadronic interactions at higher energies. For hadronic interactions at the lowest energy range (below 70 MeV) and nuclear deexcitation coupled with the higher-energy QGS model, GEANT4’s precompound (P) model is employed. Additionally, the Bertini Cascade (BERT) model is applied for hadronic interactions at the lower energy range. The High-Precision neutron (HP) model is used for neutron elastic and inelastic interactions below 20 MeV, ensuring high accuracy in modeling these interactions. The detector simulation incorporates the decay processes of unstable residual nuclei, utilizing information such as decay types, endpoints, and lifetimes sourced from the nuclear database [49].
To be concrete, we take the JUNO detector [17] as our reference. For simplicity, the detector simulation excludes the optical process, and the quenching effect in the LS is accounted for using Birk’s equation described in Ref. [50]. Consequently, the detector simulation converts the kinetic energies of final-state particles from NC interactions into quenching energy, representing the visible energy. Relevant time and vertex information is recorded, allowing the selection of events of interest based on associated information, such as single signals and coincident events. This work specifically focuses on NC backgrounds within the visible energy range below 100 MeV.
3 Characteristics of -C NC Interactions
Using the -C NC interaction prediction method detailed in Sec. 2 and adopting representative models from Tab. 1, we have simulated the -C NC interaction in the LS detector. This involves the simulation of secondary interactions caused by final-state particles. The section delves into the characteristics of these interactions across simulation stages, covering -nucleon NC interactions, FSI, deexcitation of the residual nucleus, and secondary interactions of final-state particles in the detector. We analyze the model variations concerning the -nucleon NC interactions and final-state information, investigating the impact of FSI, nuclei deexcitation, and secondary interactions on the final-state information.
@@ 3 Characteristics of -C NC Interactions
First, let us consider the -nucleon NC interactions, which are fundamental to -C NC interactions. Understanding these interactions is crucial for comprehending the dynamics of neutrinos and their impact on nucleons within C nuclei. This understanding directly influences the probability of NC interaction between neutrinos and C, determining corresponding event rates. Due to a lack of data constraints, there exists significant uncertainty in this process. Representative models in Tab. 1 aid in understanding average cross-sections between and nucleons within C, as well as the event rate distributions of -C NC interactions. To understand the model variations of these models in low and high initial neutrino energy () ranges, we divide the NC interactions into two groups:
@@ 3.1.1 Cross-sections
Fig. 2 illustrates the average cross-sections between and nucleons within C, based on the representative models listed in Tab. 1. The top and bottom panels depict the cross-sections for neutrinos and anti-neutrinos, respectively. The three panels, from left to right, represent the inclusive cross-sections, exclusive cross-sections in group I, and exclusive cross-sections in group II, respectively. In all panels, the shaded bands show the variations of across the range of values from the old models, as a function of the initial neutrino energy. The smallest cross-section values in all panels are obtained from NuWro, Model-N5, which uses GeV for QE and the SF nuclear model. On the other hand, the maximum values predominantly come from NuWro, Model-N3, and Model-N4. Model-N3 employs GeV for the QE process, and Model-N4 uses the TEM model to predict the contribution of the process, resulting in a larger cross-section below 1 GeV, but the TEM effect is absent in the antineutrino sector. Note that in group II, except for Model-N5, the other old models show quite similar cross-section results.
@@
The cross-section results from the new models added in this work are depicted using lines with distinct colors and styles. Compared to Model-N4 and Model-N3, these new models exhibit lower cross-sections. Note that the cross-sections concern the initial neutrino interactions so that the cross-sections between Model-G2 and Model-G3 are the same. Model-N7 with the SF nuclear model still offers the smallest value of the inclusive cross-section, consistent with Model-N5 using a different NuWro version.
@@
In this work, we specifically investigate model variations for the cross-sections in the low energy range, particularly focusing on the QE and interactions. In addition to the insights gained from variations in inclusive cross-sections, we also observed significant differences when examining the specificities of cross-sections in QE and interactions, as depicted in the middle panel of Fig. 2. These insights are based on the new models used for illustration. Firstly, the nuclear model has a substantial impact on the cross-section at low energies. For instance, it can be observed by comparing Model-G2 and Model-G4 in GENIE with the same settings but different nuclear models (LFG for Model-G2 and BRRFG for Model-G4) and comparing Model-N6 and Model-N7 in NuWro with the same settings but different nuclear models (LFG for Model-N6 and SF for Model-N7). Secondly, consistent alignment of cross-sections from Model-G2 and Model-N6 takes place above 300 MeV. However, significant deviations are observed at lower energies. This discrepancy is likely attributed to different treatments of the initial nuclear state in GENIE and NuWro. Thirdly, precise model descriptions of processes are essential for accurate predictions of atmospheric neutrino NC interactions at low energies. These variations of cross-sections regarding the initial interactions from all representative models are considered as a systematic uncertainty in our calculations.
3.1.2 -C NC interaction rates
@@
We utilize the atmospheric (anti)neutrino fluxes at the JUNO site and the cross-section distributions from Fig. 2 to calculate the event rate of -C NC interactions. The relevant energy range of the atmospheric neutrinos used in this work is 100 MeV 10 GeV, and the total number of the target C nuclei per kiloton LS is around . The event rates of group I and group II of the atmospheric neutrinos with C nuclei are presented in Fig. 3 by summing up all the neutrino species. The rates are depicted concerning the initial neutrino energy and the energy transfer, shown in the left and right panels, respectively. The energy transfer () is defined as the energy difference between incoming and outgoing neutrinos.
@@ 3.2 Final-state Information
Then, proceeding to the examination of final-state particles in NC events becomes crucial for effective tagging. In -nucleon NC interactions, neutrinos depart from the nuclei, transferring energy to nucleons as depicted in the right panel of Fig. 3. In the scenario where nucleons are free, the final-state of the QE process comprises solely single nucleon productions. However, these nucleons are bound within the C nuclei. Subsequent processes such as FSI and nucleus deexcitation have the potential to expel additional nucleons, resulting in the emission of multiple final-state particles, including , , , and others. To investigate the influence of FSI and deexcitation on final-state particle production, we employ new models to analyze and compare their effects across different FSI models, evaluating outcomes when deexcitation is either activated or deactivated.
@@ 3.2.1 Production rates of final-state particles
In the new GENIE models, Model-G2 exhibits the lowest rates, while Model-G4 shows the highest rates. The distinction between these models lies in the choice of both nuclear and FSI models. Specifically, Model-G4, employing the BRRFG model, produces around 8% higher rate of neutrons compared to Model-G2 with the LFG model. Additionally, Model-G3, featuring the hA2018 FSI model, yields about 6% higher rates of neutrons compared to Model-G2 with the hN2018 FSI model.
When excluding nucleons with zero kinetic energies in the final-state, Model-G2’/4’ with the hN2018 FSI model exhibit a 25% lower rate of neutrons compared to the associated models. Similarly, Model-G3’ with the hA2018 FSI model shows a 10% lower rate of neutrons in comparison. Therefore, in the new GENIE’ models, Model-G2’ still exhibits the lowest rates, while Model-G3’ shows the highest rates.
In the new NuWro model variations, Model-N6 with the LFG model generates 18% higher neutron rates compared to Model-N7 with the SF nuclear model. Furthermore, NuWro models exhibit significantly lower neutron rates than those from GENIE models. These differences may be subject to constraints imposed by future measurements of -C NC interactions using large LS detectors.
When comparing production rates with and without deexcitation, neutron production rates, including deexcitation, show an enhancement of around (13 19)% compared to those without deexcitation for all new models. Additionally, deexcitation induces almost all ’s, ’s, ’s, He’s, and ’s particles. Note that variations in the models of products from deexcitation are associated with the model variations of the excited residual nucleus.
@@ 3.2.2 Neutron multiplicity
In Fig. 6, we examine the neutron multiplicity in the final-state of NC interactions involving -C. Considering that the predominant neutron multiplicities are below three, the event rates for neutron multiplicities less than three are highlighted in the subpanels of the upper panels of Fig. 6. Due to the significant similarity in distributions between Model-G2 and Model-G4, as well as between Model-G2’ and Model-G4’, Model-G4 and Model-G4’ are excluded from the illustrations. Excluding neutrons from the deexcitation process, the events with neutron multiplicities less than three account for approximately 91% in the new GENIE models using the hN2018 FSI model (i.e., Model-G2 and Model-G4) and hA2018 FSI model (i.e., Model-G3). This calculation includes neutrons with zero kinetic energies in the final-state. If we exclude these neutrons in Model-G 2’, 3’, 4’, the events with neutron multiplicities less than three increase to 95% in Model-G2’ or Model-G4’, while the fraction remains unchanged in Model-G3’. This difference arises because neutrons with zero kinetic energies from the hN2018 FSI model impact all neutron multiplicities, whereas those from the hA2018 FSI model only affect neutron multiplicities less than three. Compared to GENIE models, NuWro models produce around 99% events with neutron multiplicities less than three. For comparison purposes, the middle panels of Fig. 6 display the ratio of results related to multiplicities less than three to those in Model-G2. The neutrons induced by the deexcitation process increase the multiplicity. The parameter, , defined as the ratio of event rates for NC events with deexcitation to those for NC events without deexcitation, concerning multiplicities less than three is illustrated in the lowest panel of Fig. 6.
@@
Fig. 7 presents a summary of the final event rates for QE and interactions of -C as a function of tagged neutron multiplicity, where tagged neutron multiplicity is defined as the numbers of neutrons captured on H or C atoms in the LS after an NC interaction. The notations and patterns of the histograms precisely mirror those in Fig. 6. Specifically, the parameter is defined as the ratio of event rates for NC events with the deexcitation process to those for NC without the deexcitation process, relative to tagged neutron multiplicities less than three. We expect that the event rates linked to tagged neutron multiplicities in the LS detector should align with the event rates of produced neutron multiplicities as depicted in Fig. 6. However, what we observe are those after secondary interactions in LS. These are phenomena that occur when final-state particles, initially produced by a primary interaction, subsequently interact within the detector. For instance, the energetic fast neutrons from the -C NC interactions may disappear or produce other neutrons due to their inelastic interactions with C. This effect has been investigated by using TALYS to calculate the cross-sections of the exclusive reaction channels at different incident neutron energies [32, 51]. In the case of fast neutrons with kinetic energies below 10 MeV, the primary interaction between neutrons and C is elastic scattering (C). However, as the energy surpasses 10 MeV, the inelastic scattering cross-section of C experiences a substantial increase, leading to competition with the elastic scattering process. Comparing Fig.6 and Fig.7, it is evident that the event rates for neutron multiplicities equal to one decrease significantly due to the secondary interaction effects. Furthermore, the differences between models are diminished as a result of the secondary interaction effects.
@@
Focusing on cases with only one neutron production in the final-state, the event rate for NC events with a neutron multiplicity equal to one in Fig. 6 is categorized by exclusive channels, as depicted in the top panel of Fig. 8. This panel highlights the predominant influence of processes leading to the production of C and B in NC events without deexcitation. Notably, there is a significant increase in event rates for Be, Li, B, Be, and Li, as well as the “others” category in NC events with deexcitation. Simultaneously, event rates associated with C and B processes experience substantial reduction due to the further knockout of one or more nucleons during deexcitation.
@@
Upon scrutinizing model variations, the maximum difference is observed in NC events associated with B for the new GENIE model variations. However, considering that nucleons with zero kinetic energy remain within the nuclei, the disparity in NC events with B is significantly mitigated in new GENIE’ model variations, resulting in lower event rates. Meanwhile, the event rate for NC events with C sees a noteworthy increase in new GENIE’ models. Regarding new NuWro model variations, the primary distinction arises in NC events with C, while other channels exhibit similarity in the absence of deexcitation scenarios. Notably, Model-N6 with the LFG nuclear model produces substantially higher event rates for NC events with C.
@@
Upon comparing Fig. 6 and Fig. 7, we observe the maximum difference exists in the event rates for neutron multiplicity equal to one. Further investigation reveals that the NC events with C are significantly decreased under the condition of only one tagged neutron in LS. This is due to the higher kinetic energies of fast neutrons from the NC events with C, making them more prone to inelastic scattering with C. This effect has been illustrated in previous work (e.g., see distributions of the kinetic energies of fast neutrons in different exclusive channels in Ref. [32] and the cross-sections of exclusive C reactions in Ref. [51]). While this is advantageous for reducing NC background levels in related topics [51], it poses a challenge for measurements of NC background based on the triple-coincident signatures of the NC events with C [32].

### v2

@@ 2.3 GEANT4-based Detector Simulation
GEANT4 (version 4.10.p02) is utilized to simulate the propagation of final-state particles resulting from -12C NC interactions within the LS detector. The simulation employs the QGSPBERT physics list, with the Quark Gluon String (QGS) model applied for hadronic interactions at higher energies. For hadronic interactions at the lowest energy range (below 70 MeV) and nuclear deexcitation coupled with the higher-energy QGS model, GEANT4’s precompound (P) model is employed. Additionally, the Bertini Cascade (BERT) model is applied for hadronic interactions at the lower energy range. The High-Precision neutron (HP) model is used for neutron elastic and inelastic interactions below 20 MeV, ensuring high accuracy in modeling these interactions. The detector simulation incorporates the decay processes of unstable residual nuclei, utilizing information such as decay types, endpoints, and lifetimes sourced from the nuclear database [56].
To be specific, we use the JUNO detector [18] as our reference. For simplicity, the detector simulation excludes the optical process, and the quenching effect in the LS is accounted for using Birk’s equation described in Ref. [57]. Consequently, the detector simulation converts the kinetic energies of final-state particles from NC interactions into quenching energy, representing the visible energy. Relevant time and vertex information is recorded, allowing the selection of events of interest based on associated information, such as single signals and coincident events. This work is specifically dedicated to investigating NC backgrounds within the visible energy range below 100 MeV.
3 Characteristics of -12C NC Interactions
Using the -12C NC interaction prediction method detailed in Sec. 2 and adopting representative models from Tab. 1, we have simulated the -12C NC interaction in the LS detector. This involves the simulation of secondary interactions caused by final-state particles. The section delves into the characteristics of these interactions across simulation stages, covering -nucleon NC interactions, FSI, deexcitation of the residual nucleus, and secondary interactions of final-state particles in the detector. We analyze the model variations concerning the -nucleon NC interactions and final-state information, investigating the impact of FSI, nuclei deexcitation, and secondary interactions on the final-state information.
@@ 3 Characteristics of -12C NC Interactions
First, let us consider the -nucleon NC interactions, which are fundamental to -12C NC interactions. Understanding these interactions is crucial for comprehending the dynamics of neutrinos and their impact on nucleons within 12C nuclei. This understanding directly influences the probability of NC interaction between neutrinos and 12C, determining corresponding event rates. Due to a lack of data constraints, there is a notable level of uncertainty associated with this process. Representative models in Tab. 1 aid in understanding average cross-sections between and nucleons within 12C, as well as the event rate distributions of -12C NC interactions. To understand the model variations of these models in low and high initial neutrino energy () ranges, we divide the NC interactions into two groups:
@@ 3.1.1 Cross-sections
Fig. 2 illustrates the average cross-sections between and nucleons within 12C, based on the representative models listed in Tab. 1. The top and bottom panels depict the cross-sections for neutrinos and anti-neutrinos, respectively. The three panels, from left to right, represent the inclusive cross-sections, exclusive cross-sections in group I, and exclusive cross-sections in group II, respectively. In all panels, shaded bands show the variations of across the range of values from old models, as a function of the initial neutrino energy. The smallest cross-section values in all panels are obtained from NuWro, Model-N5 ( GeV for QE and SF). On the other hand, the maximum values predominantly come from NuWro, Model-N3/4. Model-N3 employs GeV for the QE process, and Model-N4 uses the TEM model to predict the contribution of the process, resulting in a larger cross-section below 1 GeV, but the TEM effect is absent in the antineutrino sector. Note that in group II, except for Model-N5, the other old models show quite similar cross-section results.
@@
The cross-section results from the new models added in this work are depicted using lines with distinct colors and styles. Compared to Model-N3/4, these new models exhibit lower cross-sections. Note that the cross-sections concern the initial neutrino interactions so that the cross-sections between Model-G2 and Model-G3 are the same. Model-N7 (SF) still offers the smallest value of the inclusive cross-section, consistent with Model-N5 using a different NuWro version.
@@
In this work, we specifically investigate model variations for the cross-sections in the low energy range, particularly focusing on the QE and interactions. In addition to the insights gained from variations in inclusive cross-sections, we also observed significant differences when examining the specificities of cross-sections in QE and interactions, as depicted in the middle panel of Fig. 2. These insights are based on the new models used for illustration. Firstly, the nuclear model significantly influences the cross-section at low energies. This impact can be discerned by comparing models with identical settings but different nuclear models. For instance, a comparison between Model-G2 (LFG) and Model-G4 (BRRFG), or between Model-N6 (LFG) and Model-N7 (SF), clearly illustrates this nuclear model effect. Secondly, consistent alignment of cross-sections from Model-G2 and Model-N6 takes place above 300 MeV. However, significant deviations are observed at lower energies. This discrepancy is likely attributed to different treatments of the initial nuclear state in GENIE and NuWro. Thirdly, precise model descriptions of processes are essential for accurate predictions of atmospheric neutrino NC interactions at low energies. These variations of cross-sections regarding the initial interactions from all representative models are considered as a systematic uncertainty in our calculations.
3.1.2 -12C NC interaction rates
@@
We utilize the atmospheric (anti)neutrino fluxes at the JUNO site and the cross-section distributions from Fig. 2 to calculate the event rate of -12C NC interactions. The relevant energy range of the atmospheric neutrinos used in this work is 100 MeV 10 GeV, and the total number of the target 12C nuclei per kiloton LS is around . The event rates of group I and group II of the atmospheric neutrinos with 12C nuclei are presented in Fig. 3 by summing up all the neutrino species. The rates are depicted concerning the initial neutrino energy and the energy transfer, shown in the left and right panels, respectively. The energy transfer () is defined as the energy difference between incoming and outgoing neutrinos.
@@ 3.2 Final-state Information
Then, proceeding to the examination of final-state particles in NC events becomes crucial for effective tagging. In -nucleon NC interactions, neutrinos depart from the nuclei, transferring energy to nucleons as depicted in the right panel of Fig. 3. In the scenario where nucleons are free, the final-state of the QE process comprises solely single nucleon productions. However, these nucleons are bound within the 12C nuclei. Subsequent processes such as FSI and nucleus deexcitation have the potential to expel additional nucleons, resulting in the emission of multiple final-state particles, including , , , and others. To investigate the influence of FSI and deexcitation on final-state particle production, we employ new models to analyze and compare their effects across different FSI models, evaluating outcomes when deexcitation is either activated or deactivated.
@@ 3.2.1 Production rates of final-state particles
In the new GENIE models, Model-G2 exhibits the lowest rates, while Model-G4 shows the highest rates. The distinction between these models lies in the choice of both nuclear and FSI models. Specifically, Model-G4 (BRRFG), produces around 8% higher rate of neutrons compared to Model-G2 (LFG). Additionally, Model-G3 (A), yields about 6% higher rates of neutrons compared to Model-G2 (N).
When excluding nucleons with zero kinetic energies in the final-state, Model-G2’/4’ (N) exhibit a 25% lower rate of neutrons compared to the associated models. Similarly, Model-G3’ (A) shows a 10% lower rate of neutrons in comparison. Therefore, in the new GENIE’ models, Model-G2’ still exhibits the lowest rates, while Model-G3’ shows the highest rates.
In the new NuWro model variations, Model-N6 (LFG) generates 18% higher neutron rates compared to Model-N7 (SF). Furthermore, NuWro models exhibit significantly lower neutron rates than those from GENIE models. These differences may be subject to constraints imposed by future measurements of -12C NC interactions using large LS detectors.
When comparing production rates with and without deexcitation, neutron production rates, including deexcitation, show an enhancement of around (13 19)% compared to those without deexcitation for all new models. Additionally, deexcitation induces almost all ’s, ’s, ’s, 3He’s, and ’s particles. Note that variations in the models of products from deexcitation are associated with the model variations of the excited residual nucleus.
@@ 3.2.2 Neutron multiplicity
In Fig. 6, we examine the neutron multiplicity in the final-state of NC interactions involving -12C. Considering that the predominant neutron multiplicities are below three, the event rates for neutron multiplicities less than three are highlighted in the subpanels of the upper panels of Fig. 6. Due to the significant similarity in distributions between Model-G2 and Model-G4, as well as between Model-G2’ and Model-G4’, Model-G4/4’ is excluded from the illustrations. Excluding neutrons from the deexcitation process, the events with neutron multiplicities less than three account for approximately 91% in the new GENIE models using the N FSI model (i.e., Model-G2/4) and A FSI model (i.e., Model-G3). This calculation includes neutrons with zero kinetic energies in the final-state. If we exclude these neutrons in Model-G 2’, 3’, 4’, the events with neutron multiplicities less than three increase to 95% in Model-G2’/4’, while the fraction remains unchanged in Model-G3’. This difference arises because neutrons with zero kinetic energies from the N FSI model impact all neutron multiplicities, whereas those from the A FSI model only affect neutron multiplicities less than three. Compared to GENIE models, NuWro models produce around 99% events with neutron multiplicities less than three. For comparison purposes, the middle panels of Fig. 6 display the ratio of results related to multiplicities less than three to those in Model-G2. The neutrons induced by the deexcitation process increase the multiplicity. The parameter, , defined as the ratio of event rates for NC events with deexcitation to those for NC events without deexcitation, concerning multiplicities less than three is illustrated in the lowest panel of Fig. 6.
@@
Fig. 7 presents a summary of the final event rates for QE and interactions of -12C as a function of tagged neutron multiplicity, where tagged neutron multiplicity is defined as the numbers of neutrons captured on H or C atoms in the LS after an NC interaction. The notations and patterns of the histograms precisely mirror those in Fig. 6. Specifically, the parameter is defined as the ratio of event rates for NC events with the deexcitation process to those for NC without the deexcitation process, relative to tagged neutron multiplicities less than three. We expect that the event rates linked to tagged neutron multiplicities in the LS detector should align with the event rates of produced neutron multiplicities as depicted in Fig. 6. However, what we observe are those after secondary interactions in LS. For instance, the energetic fast neutrons from the -12C NC interactions may disappear or produce other neutrons due to their inelastic interactions with 12C. This effect has been investigated by using TALYS to calculate the cross-sections of the exclusive reaction channels at different incident neutron energies [45, 58]. In the case of fast neutrons with kinetic energies below 10 MeV, the primary interaction between neutrons and 12C is elastic scattering (C). However, as the energy surpasses 10 MeV, the inelastic scattering cross-section of C experiences a substantial increase, leading to competition with the elastic scattering process. Comparing Fig.6 and Fig.7, it is evident that the event rates for neutron multiplicities equal to one decrease significantly due to the secondary interaction effects. Furthermore, the differences between models are diminished as a result of the secondary interaction effects.
@@
Focusing on cases with only one neutron production in the final-state, the event rate for NC events with a neutron multiplicity equal to one in Fig. 6 is categorized by exclusive channels, as depicted in the top panel of Fig. 8. This panel highlights the predominant influence of processes leading to the production of 11C and 10B in NC events without deexcitation. Notably, there is a significant increase in event rates for 9Be, 6Li, 8B, 7Be, and 7Li, as well as the “others” category in NC events with deexcitation. Simultaneously, event rates associated with 11C and 10B processes experience substantial reduction due to the further knockout of one or more nucleons during deexcitation.
@@
Upon scrutinizing model variations, the maximum difference is observed in NC events associated with 10B for the new GENIE model variations. However, considering that nucleons with zero kinetic energy remain within the nuclei, the disparity in NC events with 10B is significantly mitigated in new GENIE’ model variations, resulting in lower event rates. Meanwhile, the event rate for NC events with 11C sees a noteworthy increase in new GENIE’ models. Regarding new NuWro model variations, the primary distinction arises in NC events with 11C, while other channels exhibit similarity in the absence of deexcitation scenarios. Notably, Model-N6 with the LFG nuclear model produces substantially higher event rates for NC events with 11C.
@@
Upon comparing Fig. 6 and Fig. 7, we observe the maximum difference exists in the event rates for neutron multiplicity equal to one. Further investigation reveals that the NC events with 11C are significantly decreased under the condition of only one tagged neutron in LS. This is due to the higher kinetic energies of fast neutrons from the NC events with 11C, making them more prone to inelastic scattering with 12C. This effect has been illustrated in previous work (e.g., see distributions of the kinetic energies of fast neutrons in different exclusive channels in Ref. [45] and the cross-sections of exclusive C reactions in Ref. [58]). While this is advantageous for reducing NC background levels in related topics [58], it poses a challenge for measurements of NC background based on the triple-coincident signatures of the NC events with 11C [45].